# Robust Adaptation of Large Multimodal Models for Retrieval Augmented Hateful Meme Detection

### v3

@@ 4.9 Comparing Rationales Generated by Models
We compare meme explanations generated by the SFT and RA-HMD fine-tuned Qwen2-VL-7B models (Table 1 rows 15 and 16). RA-HMD is evaluated after Stage-1 fine-tuning, as Stage-2 does not fine-tune the Language Model Head for language generation.
Following prior work Yang et al. (2023), we assess explanation quality using an LLM judge using pair-wise comparison and also a rubric-based evaluation. We evaluate the SFT and RA-HMD fine-tuned Qwen2-VL-7B models on the validation set of HatefulMemes dataset where the ground truth human annotated reasoning is available as a reference Hee et al. (2023). The LLM judge measures how closely model-generated explanations align with human rationales. Full details of the evaluation procedure are provided in Appendix L.
@@
These results are visualized in Figure 1.
For the rubric-based evaluation (scored on a scale of 0–10), the SFT baseline achieved an average score of 6.6, while RA-HMD obtained a higher score of 7.5, further indicating stronger alignment with human rationales.
Based on the analysis of the generated explanations, we find that improvements in classification accuracy are supported by a deeper semantic understanding of memes. For challenging examples, where comprehension of background events or fine-grained details from the image is required, the RA–HMD–tuned system generates explanations that are both more accurate and more informative.
@@ 5 Conclusion
Hate speech is described using various terminologies, including online harassment, online aggression, cyberbullying, and harmful speech. The United Nations Strategy and Plan of Action on Hate Speech acknowledges that definitions of hate speech can be controversial and subject to debate Nderitu (2020). Similarly, the UK Online Harms White Paper highlights that certain harms may be insufficiently defined Woodhouse (2022).
We acknowledge that the definition of hate speech can be subjective and varies across different cultural and legal contexts. To this end, we evaluate our methods on six widely used meme classification datasets, allowing for generalization across different definitions of hate speech. As the discourse on defining hate speech evolves, we align our research with this ongoing process and plan to incorporate new datasets as they become available.
In our error analysis, we find that the system is unable to recognize subtle visual details in memes. Enhancing image understanding through a more powerful vision encoder could further improve performance, which we leave for future work.
@@
Detailed experimental setups, implementation specifics, and hyperparameter settings are provided in Appendix B to ensure reproducibility. The source code will be released upon publication.
@@ Appendix B Experiment Setup and Implementation Details
We conducted our experiments on a workstation equipped with NVIDIA RTX 3090 was used for the experiments.
@@ B.1 LLaVA and Qwen2-VL experiments
For Qwen2-VL fine-tuning, we employ the officially recommended fine-tuning library LLaMA-Factory 0.9.1444https://github.com/hiyouga/LLaMA-Factory with official hyperparameter settings for downstream tasks in both the SFT and RA-HMD stage 1 fine-tuning. For few-shot learning with Qwen2-VL, we follow the official multi-round conversation prompt format to ensure consistency with the model’s intended usage.
@@ B.2 Hyperparameters for MLP and Stage 2 Fine-tuning
RGCL hyperparameter Value
@@ Appendix I Evaluation on General Vision-Language Benchmarks
[Instruction]
Please act as an impartial judge and evaluate the quality of the responses provided by two AI assistants to the user question displayed below. You should choose the assistant that follows the users instructions and answers the users question more accurately.
When choosing the assistant, please consider the true answers below:
Target: {true_label}{reference_info}
Implied Statement: This is a hateful meme with text: "{meme_text}"
Your evaluation should consider which response is more similar to the true answers. Begin your evaluation by comparing the two responses and provide a short explanation. Avoid any positional biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Do not favor certain names of the assistants. Be as objective as possible. After providing your explanation, output your final verdict by strictly following this format: "[[A]]" if assistant A is more accurate, "[[B]]" if assistant B is more accurate, and "[[C]]" for a tie.
[Question]
Determine whether the following post is offensive or not. You must choose one of the options.
Post: A meme with text: "{meme_text}"
Options:
(A) Offensive
(B) Not offensive
@@
Additionally, we conduct a rubric-based analysis, following the procedure in Yang et al. 2023. Specifically, we prompted the LLM judge to assign a score from 1 to 10 to each system’s rationale, based on how well it aligns with the human-written explanations. We assess the SFT and RA-HMD tuned Qwen2-VL-7B on the dev-seen split of the Facebook HatefulMemes dataset. Our results show that the SFT baseline received an average score of 6.6, while RA-HMD scored higher at 7.5, further indicating stronger alignment with human rationales.

### v4

@@ 4.9 Comparing Rationales Generated by Models
We compare meme explanations generated by the SFT and RA-HMD fine-tuned Qwen2-VL-7B models (Table 1 rows 15 and 16) on the validation set of the HatefulMemes dataset, where human-annotated rationales are available as ground-truth references Hee et al. (2023). RA-HMD is evaluated after Stage-1 fine-tuning, since Stage-2 does not fine-tune the Language Model Head for language generation.
Following prior work Yang et al. (2023), we evaluate explanation quality using two approaches based on LLM-as-judge:
  • pair-wise comparison,
  • rubric-based evaluation.
The LLM judge measures how closely model-generated explanations align with human rationales. Full details of the evaluation protocol are provided in Appendix L.
@@
and these results are visualized in Figure 1.
For the rubric-based evaluation (scored on a scale of 0–10), the SFT baseline achieved an average score of 4.9, while RA-HMD obtained a higher score of 5.6, further indicating stronger alignment with human rationales.
Based on the analysis of the generated explanations, we find that improvements in classification accuracy are supported by a deeper semantic understanding of memes. For challenging examples, where comprehension of background events or fine-grained details from the image is required, the RA–HMD fine-tuned system generates explanations that are both more accurate and more informative. Nevertheless, with an average score of 5.6, the performance remains far from perfect. Further discussion of this aspect is provided in Limitations.
@@ 5 Conclusion
Insufficient Definition of Hate Speech: Hate speech is described using various terminologies, including online harassment, online aggression, cyberbullying, and harmful speech. The United Nations Strategy and Plan of Action on Hate Speech acknowledges that definitions of hate speech can be controversial and subject to debate Nderitu (2020). Similarly, the UK Online Harms White Paper highlights that certain harms may be insufficiently defined Woodhouse (2022).
Variation in the Definition of Hate Speech: We acknowledge that the definition of hate speech can be subjective and varies across different cultural and legal contexts. To this end, we evaluate our methods on six widely used meme classification datasets, allowing for generalization across different definitions of hate speech. As the discourse on defining hate speech evolves, we align our research with this ongoing process and plan to incorporate new datasets as they become available.
Fine-grained Vision Understanding: In our error analysis, we find that the system is unable to recognize subtle visual details in memes. Enhancing image understanding through a more powerful vision encoder could further improve performance, which we leave for future work.
Preference and RL-based Tuning Methods for Reasoning: In this work, our analysis and methodology primarily build on SFT, which may partly explain the limitations in reasoning quality. Future research could benefit from integrating reasoning capabilities from emerging models such as OpenAI-o1 OpenAI (2024b) and DeepSeek-R1 DeepSeek-AI (2025). A deeper investigation into preference-based and RL-based tuning methods, including DPO and GRPO, may prove valuable for incentivizing stronger reasoning abilities in LMMs.
@@
Detailed experimental setups, implementation specifics, and hyperparameter settings are provided in Appendix B to ensure reproducibility. The source code can be accessed from GitHub.
@@ Appendix B Experiment Setup and Implementation Details
We conducted our experiments on a workstation equipped with an NVIDIA RTX 3090.
@@ B.1 LLaVA and Qwen2-VL experiments
For Qwen2-VL fine-tuning, we employ the officially recommended fine-tuning library LLaMA-Factory 0.9.1444https://github.com/hiyouga/LLaMA-Factory with official hyperparameter settings for downstream tasks in both the SFT and RA-HMD stage 1 fine-tuning. The only modifications are the LoRA hyperparameters: we employ a larger rank (64) and alpha (128), which are kept fixed throughout all experiments. We fix these LoRA hyperparameters throughout.
Note that in the open-source version on GitHub  [^5] https://github.com/JingbiaoMei/RGCL, we update the LLaMA-Factory version to support Qwen2.5-VL fine-tuning for later study, but the reported results in this paper are based on the version noted above. We further provide scripts to run Qwen2.5-VL-7B and Qwen2.5-VL-32B experiments. Surprisingly, both these two models achieve almost identical results compared to the Qwen2-VL-7B model reported in this paper. Given that, we do not further scale up the parameter to 72B for this study.
For few-shot learning with Qwen2-VL, we follow the official multi-round conversation prompt format to ensure consistency with the model’s intended usage.
@@ B.2 Hyperparameters for MLP and Stage 2 Fine-tuning
Stage 2 Contrastive Learning Hyperparameter Value
@@ Appendix I Evaluation on General Vision-Language Benchmarks
Please act as an impartial judge and evaluate the quality of the model-generated reasoning provided by two AI assistants. You will compare the two model-generated reasoning with the reference human reasoning for a hateful meme.
Your evaluation should consider which response is more similar to the true answers. Begin your evaluation by comparing the two responses and provide a short explanation. After providing your explanation, output your final verdict by strictly following this format: "[[A]]" if assistant A is more accurate, "[[B]]" if assistant B is more accurate, and "[[C]]" for a tie.
@@
[The Start of Reference Answer]
{reference}
[The End of Reference Answer]
@@
Additionally, we conduct a rubric-based analysis, following the procedure in Yang et al. 2023. Specifically, we prompted the LLM judge to assign a score from 0 to 10 to each system’s rationale, based on how well it aligns with the human-written explanations.
For reference, the prompt we use is:
Compare the model-generated reasoning with the reference human reasoning for this hateful meme.
Reference: {reference_reasoning}
Model: {model_reasoning}
Rate how well the model reasoning aligns with the reference on a scale of 0-10:
- 9-10: Excellent alignment, captures all key points
- 7-8: Good alignment, captures most key points
- 5-6: Satisfactory alignment, captures some key points
- 3-4: Poor alignment, misses many key points
- 1-2: Very poor alignment, minimal understanding
- 0: Completely wrong or unrelated
Score: [0-10]
Explanation: [1-2 sentences]
We assess the SFT and RA-HMD tuned Qwen2-VL-7B on the dev-seen split of the Facebook HatefulMemes dataset. Our results show that the SFT baseline received an average score of 4.9, while RA-HMD scored higher at 5.6, further indicating stronger alignment with human rationales.